OpenStudy (anonymous):

Medal, Fan, and Testimonial!!!!!!!!!!!!! HHHHHEEEEEELLLLLLLPPPPPPP!!!!!!! Solve the equation and identify any extraneous solutions

OpenStudy (anonymous):

\[2\sqrt{x} = -4\]

OpenStudy (anonymous):

@damonte212

OpenStudy (anonymous):

let me do that equation again 

OpenStudy (anonymous):

@BloomLocke367

OpenStudy (anonymous):

@mathtastic101

OpenStudy (anonymous):

\[\sqrt{-2i}\] Just a thought. Don't take it as completely correct

OpenStudy (anonymous):

Can you show me how you got to that?

OpenStudy (anonymous):

@lollololoool

OpenStudy (anonymous):

@damonte212 @iloveyouboo @BloomLocke367

OpenStudy (anonymous):

@Kentekai

OpenStudy (anonymous):

I cheated and used a calculator hope this helps or is correct and stuff

OpenStudy (anonymous):

Divide the 2 on both sides and you'll have \[\sqrt{x}=-2\] then square both sides to remove the radical and you'll end up with \[x=4\] My apologies

OpenStudy (anonymous):

so what are the extraneous?

OpenStudy (anonymous):

And that cannot be right anyway. it would be 2/x (that is the square root sign) -4. so it would be /x = 8 right?

OpenStudy (anonymous):

Oh wait, i see what you did now. Okay, i still need extraneous solutions though.

OpenStudy (anonymous):

Yeah

OpenStudy (anonymous):

I thought it was -8 because -8/2=-4

OpenStudy (anonymous):

one moment and I'll have the answer

OpenStudy (anonymous):

Otay

OpenStudy (anonymous):

Do you mind if I close this question? We can still type and stuff, but I have a few more I need answered.

OpenStudy (anonymous):

Not a problem.